OpenStudy (anonymous):

WILL MEDAL AND FAN The formula Q = MCT where Q = heat flow, M = mass, C = specific heat, and T = change of temperature is used to calculate heat flow. Solve this formula for T. T = Q – MC T = Q + MC T = QMC T = Q divided by the quantity M times C

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (anonymous):

@IrishBoy123 could you help?

OpenStudy (irishboy123):

\(Q = MCT\) you want: \(T = something\) so you're gonna have to start dividing each side by something, right?

OpenStudy (anonymous):

ok... yes

OpenStudy (irishboy123):

so do it!! divide each side by something that gets you T = [the answer] :p

OpenStudy (anonymous):

q/t = mc/t?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

hi

OpenStudy (irishboy123):

no :( \(Q = MCT\) \(\frac{Q}{T} = \frac{MCT}{T} = \frac{MC}{1} = MC\) not that helpful, is it? you could try \(\frac{Q}{M} = \frac{MCT}{M} = \frac{CT}{1} = CT\) and do something very similar again

OpenStudy (anonymous):

ok... so I am still confused, just a whole bunch of letters

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

AHH I NEED HELP ASAP

OpenStudy (irishboy123):

" just a whole bunch of letters" the answer is supposed to be a whole bunch of letters :p and algebra is a whole bunch of letters too it's not easy so don't worry about it - but i cannot just do it for you. someone else might, or someone else might be better at explaining. i'd say go back and look at the bunch of letters and see where we got to just by doing the same thing to each side of the equals sign.

OpenStudy (irishboy123):

then finish it off

OpenStudy (anonymous):

I dont want the answer, I just need to understand it better, and I want you to help me eliminate the wrong answer choices...

OpenStudy (michele_laino):

we can start from the suggestion of @IrishBoy123 so we can write: \[\Large \frac{Q}{M} = CT\]

OpenStudy (michele_laino):

@briannam568

OpenStudy (anonymous):

i want a medal

OpenStudy (michele_laino):

now if we divide both sides of that above equation, by C, we get: \[\Large \frac{Q}{{MC}} = \frac{{CT}}{C}\] please simplify @briannam568

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do I simplify that? Ive been trying for twenty minutes? @mathstudent55

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

on the RHS, you have C on the top and bottom of the fraction. what to do?

OpenStudy (anonymous):

you cross it out, since it is canceled out, right?

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@ganeshie8

OpenStudy (johnweldon1993):

^Right \[\large \frac{Q}{MC} = \frac{\cancel{C}T}{\cancel{C}}\] So what do you have left?

OpenStudy (anonymous):

Q/MC=T!! OHH THANK YOU SOOO MUCH!

OpenStudy (irishboy123):

"you cross it out, since it is canceled out, right?" indeed, you did it well done